Find the pressure on the roof of Grand Hyatt Hotel Manila (one of the tallest building in the Philippines if its height is 1,043 ft.
1
Expert's answer
2021-10-29T05:07:15-0400

Dependency of the barometric pressure on the altitude is given by the formula


"P=P_0\exp(-\dfrac{Mg}{RT}h)"

"P_0=101.325\ kPa, g-=9.807\ m/s^2,"

"M=0.02896\ kg/mol, T=288.15\ K"

"R=8.3143\ (N\cdot m)/(mol\cdot K)"

"h=1043\ ft=1043\cdot0.3048\ m=317.9064\ m"



"P=101.325\ kPa\cdot"

"\exp(-\dfrac{0.02896\ kg/mol\cdot9.807\ m/s^2}{8.3143\ (N\cdot m)/(mol\cdot K)\cdot288.15\ K}\cdot317.9064\ m)"

"=97.775\ kPa"

"97.775\ kPa\approx28.8146\ inches\ Hg"